OpenStudy (anonymous):

James is playing in a tennis competition. He is required to play 8 matches and he must win all 8 to win the competition. He has 10 percent chance of winning the competition. In each match his probability of winning decreases by 5 percent of the preceding match's probability. What is the probability that he wins the first match?

OpenStudy (anonymous):

15.513% ???

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

14.07% ??

OpenStudy (ganpat):

10% chances for winning competition.. i.e for winning all 8 matches.. say he has probability of winning in 1st match = x% SO in every match , it reduces by 5%.. so probability for every match will be x + (x-5) + (x-10) + (x-15 ) + (x -20) + (x-25) + (x-30) + (x-35) = 10 x + 7x -140 = 10 8x = 150 x= 150/8 x = 18.75 % what say ??

OpenStudy (anonymous):

doesnt look right..

OpenStudy (anonymous):

it got to do somethig with gp's ..

OpenStudy (anonymous):

reasoning as follows put \(p\) as the probability he wins the first match, making the probability he wins the second match \(.95p\) and the third as \(.9p\) and so on since winning 8 is an AND statement, you multiply these together and set it equal to .1 you would expect the probability he wins the first first to be rather large

OpenStudy (anonymous):

looks like about .978 or around 98%

OpenStudy (jiteshmeghwal9):

GP = ??? @satellite73 sir :)

OpenStudy (anonymous):

i got it thanks. and its 90.

OpenStudy (anonymous):

really? i must have typed it in wrong

OpenStudy (anonymous):

well this is how i did it.. \[(0.95)p * (0.95)^2p *....* (0.95)^7p = 0.1\] then solved for p

OpenStudy (anonymous):

oh yeah i did it is about 92% right?

OpenStudy (anonymous):

exactly 90

OpenStudy (anonymous):

welll 0.89...

OpenStudy (anonymous):

oh wow did i make a bush league mistake sorry

OpenStudy (anonymous):

(blush)

OpenStudy (anonymous):

why 0.95, 0.9, 0.85?

OpenStudy (anonymous):

yes you have it because i am an idiot, you are exactly right

OpenStudy (anonymous):

what you have is correct, ignore me now i better get some coffee